OpenStudy (anonymous):

how do you use the slope-intercept form to write an equation of a line given its slope and y-intercept?

jimthompson5910 (jim_thompson5910):

y = mx+b is the general slope intercept form m is the slope b is the y-intercept

jimthompson5910 (jim_thompson5910):

example: y = 3x+10 is in y = mx+b form where m = 3, b = 10 the slope is m = 3 the y-intercept is b = 10

OpenStudy (jdoe0001):

how do you use the slope-intercept form to write a line written in slope-intercept form? just plug in the values then